# What You Say = What You Want? Teaching Humans to Articulate Requirements for LLMs

### v1

@@ 2.3. Requirements in Programming and Software Engineering
In software engineering and end-user software engineering, requirements describe what a human wants to achieve and how a program should behave in the world (Ko et al., 2011). They encompass all the necessary conditions, constraints, and desired outcomes to ensure the output aligns with the human’s needs and expectations. Requirement engineering stems as a field that focuses on generating and documenting requirements for software (Walia and Carver, 2009). Good quality requirements need to be accurate and complete, without commission (inclusion of irrelevant or incorrect details) and omission (exclusion of necessary details) defects (Alshazly et al., 2014). Previous studies have identified requirement engineering as a challenging skill to master. While training mechanisms have been developed to help students avoid commission and omission errors (Neto et al., 2013), a significant skill gap remains between graduates and professional engineers (Radermacher et al., 2014; Eckerdal et al., 2006).
@@
Parallels can be drawn between prompt engineering (PE) and requirements engineering. For example, adapting LLMs to diverse scenarios demands well-specified requirements, and prompt authors often need to iterate on prompts that are too ambiguous for the LLM to interpret (Petridis et al., 2024). While PE training stresses the importance of clear requirements (Meskó, 2023; Schmidt et al., 2024; Desmond and Brachman, 2024), this is often presented just as one of many prompting principles (e.g., one of 26 (Bsharat et al., 2023)), limiting its impact on prompt authors. In fact, little attention has been given to explicitly training end-users on requirement engineering in the context of prompt programs, including both natural-language-to-code and natural-language driven GPT applications. While several tools like EvalLM (Kim et al., 2023), SPADE (Shankar et al., 2024a), and EvalGen (Shankar et al., 2024b) have explored extracting user criteria to support prompt evaluation, the concept of requirements has not yet been fully emphasized in the context of prompt construction. We aim to address this gap by introducing the ROPE paradigm and developing a requirement-focused training mechanism to help end-users more effectively instruct LLMs to achieve their desired goals.
@@ 3. The ROPE Paradigm
In our work, we focus on evaluating requirements quality rather than quantity, and we operationalize requirement quality in our work (defined in Section 4.2) by adopting a taxonomy from the requirement engineering literature (Section 2.3). Ensuring requirements are accurate and complete is essential for effective LLM prompting, as poor-quality requirements can lead to harmful outcomes — e.g., missing requirement ‘‘anonymize the data’’ may make LLMs keep identifiable information in data, and vague requirement ‘‘delete harmful content’’ without a clear definition of “harmful” may cause LLMs to misinterpret sensitive topics like mental health or race as harmful.
Here we also focus on natural language requirements due to its universal understanding and tight connection to LLM prompting. While multi-modal requirements can be compiled into prompts for different LLMs, we leave this exploration for future work.
@@
We view a prompt as a super-set of requirements. It contains not only users’ customized requirements, but also other (orthogonal) factors like fluency and standard prompting tricks. Among these factors, requirements are more user-centered and LLM-agnostic — users’ goals generally remain consistent across models (e.g., GPT-4, Gemini). Fully articulating these requirements is essential, as omitted customized details can be difficult to recover automatically (explored further below). In contrast, other factors tend to be more LLM-centered: different models may prefer different writing styles, example types, or requirement order. While optimizing these factors can enhance LLM performance, they are better suited for automation rather than humans due to their (1) formulaic nature (e.g., wrapping prompts in a template consistently improves Llama-3-8b’s performance on any task (Dubey et al., 2024; lla, [n. d.])) and (2) idiosyncratic behavior (e.g., GPT-3 can be sensitive to semantically invariant edits that humans do not expect to make drastic differences (Zamfirescu-Pereira et al., 2023)). This division is crucial given the rapid evolution of LLMs, as developing a mental model for a specific LLM can be a waste of effort in the long run.
With this distinction, we view ROPE as a task-delegation strategy. In crafting executable prompt programs, users should focus on iterating and refining requirement-related components, leaving other aspects automation. We find recent advancements in prompt optimizers (Section 2.2) particularly promising. With a consistent set of user requirements, we hope future optimizers could function like program synthesizers, searching through possible “implementations” — in terms of phrasing, examples, and structure — to generate prompts that best align with each LLM’s preferences.
@@
We argue that explicitly articulating requirements becomes more crucial for what we refer to as LLM-hard tasks. These tasks require significant customization where users must intentionally deviate LLMs from their default behaviors (Qin et al., 2024; Shankar et al., 2024a; Jiang et al., 2023), making autofulfillment impossible (e.g., how many rounds of questions should be asked in Figure 2). At the time of writing, tasks that involve multi-step processes, decision-making, or context knowledge are likely more LLM-hard and thus benefit more from explicit requirements. However, this is a dynamic concept that evolves as LLMs improve.
Nowadays, LLMs are widely used by end user to create reusable prompt programs, and these prompt programs are often LLM-hard, as users frequently seek customization with various requirements, anticipating diverse inputs and expected outputs. From content creators crafting prompts for creative graphics, educators tailoring LLMs as tutors, to office workers automating workflows with GPT, people across various fields are all building natural language prompt programs. [^3] While a prompt program can output code like traditional natural language programs (Jiang et al., 2022; Raza et al., 2015), it can also produce other outputs such as videos or GPTs applications. Note that when we refer to natural language program later in the paper, we are generally describing prompt programs. We argue that all these potential prompt programmers would benefit from articulating better requirements. [^4] In contrast, more ad-hoc interactions with LLMs, such as casual conversations with the model, may require a different set of considerations, which we consider beyond the scope of this discussion.
@@ 4.1. Task Design: LLM-Hard Prompt Programs for Replication
Example & [Uncaptioned image] [Uncaptioned image]
Requirements
@@
Specificity Higher-level requirements covering diverse category, but can have subjective interpretations Lower level requirements akin to implementation specification but more verifiable
Modality Convey requirements through textual examples — users are given 2-3 chat histories as interaction examples. It is representative of status-quo prompt iteration. Convey requirements visually — users are given an interactive game. The interaction reduces ambiguity that are common in example-based representations and requires less English reading skills.
Table 1. An overview and comparison on the task types in our training.
@@
To ensure the training and evaluation process is both realistic and representative, our primary objective is to have users write natural language prompts to instruct LLMs in generating programs. Instead of focusing on open-ended tasks — which are difficult to assess and provide consistent feedback on — we aim to provide clear ground truths for evaluation and training. To this end, users are asked to write prompts that instruct LLMs to replicate a given reference program. Each task consists of two key components: (1) A gold reference program with which users can interact with to deduce requirements; and (2) A set of gold requirements that precisely reflect the program’s behavior. Ideally, an expert skilled in requirement-driven prompt engineering should be able to express all the gold requirements within their prompts, guiding the LLM to generate a program that mirrors the reference program’s behavior exactly.
This setup requires users to derive requirements from provided examples. While different from generating requirements from scratch, it captures actual challenges of requirement articulation, especially for LLM-hard tasks in real-world (as described in Section 3): Users often have access to partial inputs or examples like a demo or a user story, often need to refine their requirements by analyzing available outputs, but do not always know how to connect model outputs with abstract requirements. We discuss potential future work of open-ended tasks in Section 7.1.
@@
To cover broad types of realistic prompt programs users may write, we incorporate both GPTs applications directly powered by natural languages, as well as natural-language-to-code programs. We prioritize tasks most suitable for targeted and engaging training, through two dimensions: (1) we prioritize generic tasks over specialized ones (e.g., data science tasks), to ensure that the challenge lies in specifying requirements rather than applying domain-specific expertise, preventing construct irrelevance; and (2) we opt for tasks with visual interfaces, enabling users to interact with and visualize outputs in a manner that is accessible to novices.
With these criteria in mind, we select three GPT-driven tasks —Outline Assistant, Trip Advisor, and Email Proofreader — along with three coding tasks presented as visual game interfaces: Connect4, Tic-Tac-Toe, and Tetris. All these tasks are inspired by real-world applications and have reference requirements, though we make necessary adaptations to ensure they are LLM-hard problems. The three game tasks are sourced from the Introduction to CS course at our institution, each with reference requirements provided by the instructor. They require minimal customization, as instructors have already incorporated LLM-hard requirements to prevent students from cheating using large language models. For the three GPTs tasks, we adapt real-world GPTs prompts, [^5] Writing Assistant prompt: link, Trip Advisor prompt: link, Email Proofreader prompt: link and make them more LLM-hard by using frameworks like IFEval to introduce customized requirements such as format constraints (Zhou et al., 2023). We then derived the requirements by parsing the underlying prompts through expert annotation and pilot testing. [^6] Detailed instruction to the task, as well as the requirement breakdown are in supplemental materials.
@@
We distribute the six tasks in assessments and training session. We use Tetris and Email Proofreader in training, and the rest for pre- and post- assessments before and after training in a counterbalanced design. In pilot studies, we confirmed that (1) we could distinguish prompting novices and experts using the tasks, with the average novice performance being 28% and expert achieving full scores, (2) all tasks could be implemented by LLMs, with the average highest score being 91%, and (3) the assessment tasks were all comparably challenging for users.
@@ 4.2. Assessment Design: Requirement-Focused Intrinsic and Extrinsic Evaluation
To evaluate users’ ability to instruct LLMs effectively, we assess both the quality of the user prompts (requirements within the prompt) and the prompt’s impact on the quality of the LLM’s output. Thus, for each user task completion, we calculate the Overall Test Score as is the average of the Requirement Quality and LLM Output Quality scores:
  • Requirement Quality Score: This intrinsic metric assesses, “Does the user’s prompt accurately and comprehensively cover all requirements?” It measures the completeness and correctness of the requirements by comparing those extracted from the user’s prompt to expert-defined reference requirements for each task (described in 4.1). Drawing from the requirements defect taxonomy (Alshazly et al., 2014; Montgomery et al., 2022), we track both commission errors (incorrect, inconsistent, or ambiguous requirements) and omission errors (missing ground-truth requirements). Given the inherent ambiguity in defining requirements and the varying levels of granularity (Section 3), we break the prompt into one-sentence clauses to extract individual requirements. The requirement quality is then operationalized as the percentage of requirement clauses that are free from commission or omission defects.
  • LLM Output Quality Score: This extrinsic metric evaluates, “Can the user’s prompt successfully guide the LLM to achieve the intended goals?” It measures the proportion of desired features implemented in the LLM-generated output that align with the reference (Section 4.1). We pass users’ prompts to GPT-4o and compare the generated output to the reference program. [^7] The output is generated using a prompt similar to ‘‘Follow this prompt: {user_prompt}’’. Further details are in the supplemental materials. Since many features are difficult to evaluate automatically (e.g., GPTs behaviors cannot always be checked via rules), we rely on manual evaluation. For games, an expert interacts with the generated program to verify whether the requirement features are correctly implemented. For GPT interactions, an expert grades whether the GPTs’ replies display expected behaviors (e.g., asking a follow-up question when the user’s input is ambiguous). We grade two to three complete conversations per GPTs, with 5 turns of GPTs responses each.
@@
Three of the authors discuss to iterate the grading rubrics on the tasks during the pilot study. We confirm that expert can measure requirement quality using the percentage of correct requirement clauses in novice prompts, and that this correlates positively with the expert’s judgement (Spearman’s $\rho=0.66$). For grading in the user study (Section 5), two authors (one of whom did not participate in the rubrics development during pilot) independently grade 10% of the randomly chosen pre-post test responses. We check the inter-rater reliability between the authors’ scoring by calculating the Intraclass Correlation Coefficient (ICC) (Koo and Li, 2016), finding a strong reliability ($ICC=0.9$, 95% Confidence Interval = [0.7, 0.98]). Any discrepancies in scoring are resolved through discussion, after which the authors individually grade half of the participants’ responses. Detailed rubrics are provided in supplemental materials.
Figure 4. Our ROPE training interface, with three types of feedback on requirement defects: (A) Conversation-based hints on incomplete or inaccurate requirements, (B) Reference requirement examples to reinforce appropriately identified and expressed requirements, and (C) LLM output counterfactual to support user’s reflection on incorrect or ambiguous requirements.
@@ 4.3. Interactive Training Mechanism: Dedicated Practice and Feedback on Requirement Defects
In this training, a user (a prompt novice) is asked to replicate one Game task (Tetris) and one GPTs task (Email Proofreader), solely by describing the requirements of the given program as accurate and complete as possible. For example, to reproduce the customized Tetris game in Figure 4, the user will first start by outlining the main milestones (e.g., creating the game board, handling piece placement, etc.), and then provide more detailed specification per step (e.g., define the size of the game board). Novices are not expected to achieve full success on their first try; we use three types of requirement-focused feedback to guide users in continuously refining their requirements:
  • Textual hint and clarification, via chatbot (Figure 4A): We create a tutor chatbot to offer users a natural, conversational experience to discuss and reflect on their requirements. For instance, the chatbot may ask “What’s on top of the board?”, when a novice misses the requirement for ‘‘Tetris title rendering’’. The textual feedback encourages critical thinking on missing or incorrect requirements.
  • Reference requirement example, via requirement working document (Figure 4B): We progressively reveal reference requirements when users mention them during interaction with the chatbot. The expert-written reference examples provide reinforcement on correct requirements, helping users understand how to formalize and organize requirements.
  • LLM output counterfactual, via generated visualization (Figure 4C): For incorrect requirements that can be visually demonstrated, we generate flawed programs by implementing incorrect requirements (e.g., wrong board dimensions), or maliciously misinterpreting ambiguous requirements. For example, a vague requirement like “Use keys to move pieces” might result in a flawed Tetris game where pieces can move upward, exposing the unspecified allowed movement in requirement. The visual counterfactual is currently limited to the controllable game code generation; as GPTs outputs tend to be less predictable, we display static chat histories as illustrative examples for GPTs.
@@
We use OpenAI GPT-4o to power the interface, [^8] The state-of-the-art LLM at the time of writing: https://openai.com/index/hello-gpt-4o/. We used a temperature of 0.3 for code generation and 0.7 for other tasks. Refer to supplemental materials for our prompts. leveraging its interactive nature to adaptively provide feedback on the diverse set of possible user inputs. However, feedback generation does not rely on the model’s inherent reasoning capabilities but is anchored in our predefined reference requirements. Specifically, we always have GPT-4o to compare the current user requirements to the reference, and select the most critical defect to provide feedback on (roughly, mismatched requirements are considered most critical, then missing requirements, then ambiguous ones). Then, we have the LLM respond targetedly to the defect. For example, visual counterfactual examples on game tasks are generated by having the LLM minimally edit reference Python code based on the identified incorrect requirement. To enhance interaction, we also implement features like highlighting special variables (e.g., board, keystrokes) in conversations and as hyperlinks for easy cross-referencing across the interface.
@@
We analyze the interface log data to evaluate the feedback quality In our user study (Section 5), we collected a total of 635 interaction turns from Tetris (ranging from 11 to 73 turns per user) and 180 turns from Email Proofreader (ranging from 4 to 20 turns per user). From this, we randomly selected 10 different users and annotated conversations for 5 users per task, resulting in 151 Tetris turns, 66 Email Proofreader turns, and a total of 113 annotations per feedback type. For each LLM turn, we assess: 1) Is the feedback needed? 2) Is the feedback provided? 3) Is the feedback right? Each feedback was categorized as correct, incorrect, irrelevant, missing, or not provided. Note that the correct feedback includes both feedback that was correctly provided when needed, and feedback that was correctly not provided when unnecessary.
@@
Description of the table. Feedback Source Right Feedback Wrong Feedback IRR ($\alpha$) Correct Not Provided Total Incorrect Irrelevant Missing Total Text 69.03% 20.35% 89.38% 10.62% 0.00% 0.00% 10.62% 0.86 Requirement Doc 48.67% 38.05% 86.73% 1.77% 5.31% 6.19% 13.27% 0.88 LLM Output 3.75% 86.25% 90.00% 0.00% 5.00% 5.00% 10.00% 0.68 Total Feedback 44.44% 44.12% 88.56% 4.58% 3.27% 3.59% 11.44% 0.87

### v2

@@ 2.3. Requirements in Software Engineering and Design
In software engineering and end-user software engineering, requirement describes what a human wants to achieve and how a program should behave in the world (Ko et al., 2011). In product and engineering design, requirement is a description of the desired solution to a problem (McDonnell and Lloyd, 2014; Wang and Zeng, 2009). Requirements encompass all the necessary conditions, constraints, and desired outcomes to ensure the output aligns with the human’s needs and expectations, and specifying requirements is often a crucial step in both the design and engineering process (Dubberly, 2004). Requirements engineering stems as a field that focuses on generating and documenting requirements for software (Walia and Carver, 2009). Good quality requirements need to be accurate and complete, without commission (inclusion of irrelevant or incorrect details) and omission (exclusion of necessary details) defects (Alshazly et al., 2014). Previous studies have identified requirements engineering as a challenging skill to master. While training mechanisms have been developed to help students avoid commission and omission errors (Neto et al., 2013), a significant skill gap remains between graduates and professional engineers (Radermacher et al., 2014; Eckerdal et al., 2006).
@@
Parallels can be drawn between prompt engineering (PE) and requirements engineering. For example, adapting LLMs to diverse scenarios demands well-specified requirements, and prompt authors often need to iterate on prompts that are too ambiguous for the LLM to interpret (Petridis et al., 2024). While PE training stresses the importance of clear requirements (Meskó, 2023; Schmidt et al., 2024; Santana, 2024), this is often presented just as one of many prompting principles (e.g., one of 26 (Bsharat et al., 2023)), limiting its impact on prompt authors. In fact, little attention has been given to explicitly training end-users on requirements engineering in the context of prompt programs, including both natural-language-to-code and natural-language driven GPT applications. While several tools like EvalLM (Kim et al., 2024) and EvalGen (Shankar et al., 2024b) have explored extracting user criteria to support prompt evaluation, the concept of requirements has not yet been fully emphasized in the context of prompt construction. We aim to address this gap by introducing the ROPE paradigm and developing a requirement-focused training mechanism to help end-users more effectively instruct LLMs to achieve their desired goals.
@@ 3. The ROPE Paradigm
In our work, we focus on evaluating requirements quality rather than quantity, and we operationalize requirement quality in our work (defined in Section 4.2) by adopting a taxonomy from the requirements engineering literature (Section 2.3). Ensuring requirements are accurate and complete is essential for effective LLM prompting, as poor-quality requirements can lead to harmful outcomes — e.g., missing requirement ‘‘anonymize the data’’ may make LLMs keep identifiable information in data, and vague requirement ‘‘delete harmful content’’ without a clear definition of “harmful” may cause LLMs to misinterpret sensitive topics like mental health or race as harmful.
Here we also focus on natural language requirements due to its universal understanding and tight connection to LLM prompting. While multi-modal requirements can be compiled into prompts for different generative AI models, we leave this exploration for future work.
@@
We view a prompt as a super-set of requirements. It contains not only users’ customized requirements, but also other (orthogonal) factors like text fluency and standard prompting tricks. Among these factors, requirements are more user-centered and LLM-agnostic — users’ goals generally remain consistent across models (e.g., GPT-4, Gemini). Fully articulating these requirements is essential, as omitted customized details can be difficult to recover automatically (explored further below). In contrast, other factors tend to be more LLM-centered: different models may prefer different writing styles, example types, or requirement order. While optimizing these factors can enhance LLM performance, they are better suited for automation rather than humans due to their (1) formulaic nature (e.g., wrapping prompts in a template consistently improves Llama-3-8b’s performance on any task (Dubey et al., 2024)) and (2) idiosyncratic behavior (e.g., GPT-3 can be sensitive to semantically invariant edits that humans do not expect to make drastic differences (Zamfirescu-Pereira et al., 2023)). This division is crucial given the rapid evolution of LLMs, as developing a mental model for a specific LLM can be a waste of effort in the long run (Liu and Chilton, 2022).
With this distinction, we view ROPE as a task-delegation strategy. In crafting executable prompt programs, users should focus on iterating and refining requirement-related components, leaving other aspects automated. We find recent advancements in prompt optimizers (Section 2.2) particularly promising. With a consistent set of user requirements, we hope future optimizers could function like program synthesizers, searching through possible “implementations” — in terms of phrasing, examples, and structure — to generate prompts that best align with each LLM’s preferences.
@@
We argue that explicitly articulating requirements becomes more crucial for what we refer to as LLM-hard tasks. These tasks require significant customization where users must intentionally deviate LLMs from their default behaviors (Qin et al., 2024; Shankar et al., 2024a; Jiang et al., 2024), making autofulfillment impossible (e.g., how many rounds of questions should be asked in Figure 2). At the time of writing, tasks that involve multi-step processes, decision-making, or context knowledge are likely more LLM-hard and thus benefit more from explicit requirements. However, this is a dynamic concept that evolves as LLMs improve.
Nowadays, LLMs are widely used by end user to create reusable prompt programs, and these prompt programs are often LLM-hard, as users frequently seek customization with various requirements, anticipating diverse inputs and expected outputs. From content creators crafting prompts for creative graphics, educators tailoring LLMs as tutors, to office workers automating workflows with GPT, people across various fields are all building natural language prompt programs. [^8] While a prompt program can output code like traditional natural language programs (Jiang et al., 2022), it can also produce other outputs such as videos or GPTs applications. Note that when we refer to natural language program later in the paper, we are generally describing prompt programs. We argue that all these potential prompt programmers would benefit from articulating better requirements. [^9] In contrast, more ad-hoc interactions with LLMs, such as casual conversations with the model, may require a different set of considerations, which we consider beyond the scope of this discussion.
@@ 4.1. Task Design: LLM-Hard Prompt Programs for Replication
Table 1. An overview and comparison on the task types in our training.
@@
Example &
Requirements (Rubrics)
[Uncaptioned image] [Uncaptioned image]
@@
Specificity Higher-level requirements covering diverse categories, but can have subjective interpretations Lower level requirements akin to implementation specification but more verifiable
Modality Convey requirements through textual examples — users are given 2-3 chat histories to control for task time and interaction exposure. It is representative of status-quo prompt iteration. Convey requirements visually — users are given an interactive game. The interaction reduces ambiguity that is common with few examples and requires less English reading skills.
@@
To ensure the training and evaluation process is both realistic and representative, our primary objective is to have users write natural language prompts to instruct LLMs in generating applications. Instead of open-ended tasks — which are difficult to assess and provide consistent feedback on — we aim to provide clear ground truths for evaluation and training. To this end, users are asked to write prompts to develop an LLM app given examples.
Each task consists of two key components: (1) A set of interaction examples from which users can deduce requirements; and (2) A set of gold requirements that precisely reflect the app’s behavior. Ideally, an expert skilled in requirement-driven prompt engineering should be able to express all the gold requirements within their prompts, guiding the LLM to generate an app that produces all the interaction examples.
This setup requires users to derive requirements from provided examples, which represents a realistic form of communication from a client to a developer. It captures actual challenges of requirement articulation, especially for LLM-hard tasks in the real world (as described in Section 3): Users as LLM app developers often have access to some text inputs or visual examples like a demo or a Figma mockup, often need to refine their requirements by analyzing available outputs, but do not always know how to connect model outputs with abstract requirements. We discuss potential future work of open-ended tasks in Section 7.1.
@@
To cover broad types of realistic prompt programs users may write, we incorporate both GPTs applications directly powered by natural languages, as well as natural-language-to-code applications. We prioritize tasks most suitable for targeted and engaging training, through two dimensions: (1) we prioritize generic tasks over specialized ones (e.g., data science tasks), to ensure that the challenge lies in specifying requirements rather than applying domain-specific expertise, preventing construct irrelevance; and (2) we opt for tasks with visual interfaces, enabling users to interact with and visualize outputs in a manner that is accessible to novices.
With these criteria in mind, we select three GPT-driven tasks —Outline Assistant, Trip Advisor, and Email Proofreader — along with three coding tasks presented as visual game interfaces: Connect4, Tic-Tac-Toe, and Tetris (see examples in Table 1). The GPTs tasks provide interaction examples in terms of textual chat histories, and the Game tasks provide interaction examples via an interactive visual interface. All these tasks are inspired by real-world applications and have reference requirements, though we make necessary adaptations to ensure they are LLM-hard problems. The Game tasks are sourced from the Introduction to CS course at our institution, each with reference requirements provided by the instructor. They require minimal customization, as instructors have already incorporated LLM-hard requirements to prevent students from cheating using large language models. For the GPTs tasks, we adapt real-world GPTs prompts, [^10] Writing Assistant prompt: link, Trip Advisor prompt: link, Email Proofreader prompt: link and make them more LLM-hard by using frameworks like IFEval to introduce customized requirements such as format constraints (Zhou et al., 2023). We then derived the requirements by parsing the underlying prompts through expert annotation and pilot testing. [^11] Detailed instructions of the task, as well as the requirement breakdown, are provided in supplemental materials.
@@
We distribute the six tasks in assessments and the training session. We use Tetris and Email Proofreader in training, and the rest for pre- and post- assessments before and after training in a counterbalanced design. In pilot studies, we confirmed that (1) we could distinguish prompting novices and experts using the tasks, with the average novice performance being 28% and expert achieving full scores, (2) all tasks could be implemented by LLMs, with the average highest score being 91%, and (3) the assessment tasks were all comparably challenging for users.
@@ 4.2. Assessment Design: Requirement-Focused Intrinsic and Extrinsic Evaluation
To evaluate users’ ability to instruct LLMs effectively, we assess both the quality of the user prompts (requirements within the prompt) and the prompt’s impact on the quality of the LLM’s output. Thus, for each user task completion, we calculate its Overall Score as the average of two components, Requirement Quality and LLM Output Quality:
  • Requirement Quality: This intrinsic metric assesses, “Does the user’s prompt accurately and comprehensively cover all requirements?” It measures the correctness and completeness of the requirements by comparing those extracted from the user’s prompt to expert-defined reference requirements for each task (described in 4.1). We operationalize requirement quality as the percentage of reference requirements that are correctly described in the user’s prompt. Drawing from the requirements defect taxonomy (Alshazly et al., 2014; Montgomery et al., 2022), we mark both commission errors (incorrect, inconsistent, or ambiguous requirements) and omission errors (missing reference requirements) in users’ prompt, so a requirement is only correctly described if it is free from commission and omission defects.
  • LLM Output Quality: This extrinsic metric evaluates, “Can the user’s prompt successfully guide the LLM to achieve the intended goals?” It measures the proportion of desired features implemented in the LLM-generated output that align with the reference (Section 4.1). We pass users’ prompts to GPT-4o and compare the generated output to the reference program. [^12] The output is generated using a prompt similar to ‘‘Follow this prompt: {user_prompt}’’. Further details are in the supplemental materials. Since many features are difficult to evaluate automatically (e.g., GPTs behaviors cannot always be checked via rules), we rely on manual evaluation. For games, we interact with the generated program to verify whether the requirement features are correctly implemented. For GPT interactions, we grade whether the GPTs’ replies display expected behaviors (e.g., asking a follow-up question when the user’s input is ambiguous). We grade 10 complete conversations per GPTs.
@@
Three of the authors discuss and iterate the grading rubrics on the tasks during the pilot study. We confirm that an expert can measure requirement quality using the percentage of correct requirement clauses in novice prompts, and that this correlates positively with the expert’s judgment of overall requirement quality (Spearman’s $\rho=0.66$). For grading the Overall Score in the user study (Section 5), two authors (one of whom did not participate in the development of the rubrics during the pilot) independently grade 10% of the randomly chosen pre-post test responses. We check the inter-rater reliability between the authors’ Overall Scores by calculating the Intraclass Correlation Coefficient (ICC) (Koo and Li, 2016), finding strong reliability ($ICC=0.9$, 95% Confidence Interval = [0.7, 0.98]). Any discrepancies in scoring are resolved through discussion, after which the authors individually grade half of the participants’ responses. Detailed rubrics are provided in supplemental materials.
Figure 4. Our ROPE training interface, with three types of feedback on requirement defects: (A) Chat-based hints on incomplete or inaccurate requirements, (B) Reference requirement examples to reinforce appropriately identified and expressed requirements, and (C) LLM output counterfactual to support user’s reflection on incorrect or ambiguous requirements.
@@ 4.3. Interactive Training Mechanism: Dedicated Practice and Feedback on Requirement Defects
In this training, a user (a prompt novice) is asked to develop one Game (Tetris) and one GPTs (Email Proofreader) given the interaction examples, solely by describing the requirements of the app as accurate and complete as possible. For example, to develop the customized Tetris game in Figure 4, the user will first start by outlining the main milestones (e.g., creating the game board, handling piece placement, etc.), and then provide more detailed specification per step (e.g., define the size of the game board). Novices are not expected to achieve full success on their first try; we use three types of requirement-focused feedback to guide users in continuously refining their requirements:
  • Conversational hint and clarification, via chatbot (Figure 4A): We create a tutor chatbot to provide feedback on users’ requirement mistakes. For instance, the chatbot may ask “What’s on top of the board?”, when a user misses the requirement for ‘‘Render a title ‘Tetris’ on top of the board’’ (omission error). As shown in Figure 4A, the chatbot may suggest the user to “double-check the board dimensions” if the user incorrectly reverses the number of rows and columns for the requirement ‘‘Render the game board (8 rows by 6 columns)’’ (commission error). The textual feedback encourages critical thinking on missing or incorrect requirements and offers a natural, conversational experience to discuss and reflect on requirements.
  • Reference requirement example, via requirement working document (Figure 4B): We progressively reveal reference requirements as feedback when users correctly identify them during interaction with the chatbot. The expert-written reference examples are provided to reinforce correct requirements, helping users understand how to formalize and organize requirements.
  • LLM output counterfactual, via generated visualization (Figure 4C): If novices produce incorrect requirements that can be visually demonstrated (e.g., wrong board dimensions), we provide visual feedback by generating flawed programs that implement the incorrect requirements, or maliciously misinterpret ambiguous requirements. For example, a vague requirement like “Use keys to move pieces” might result in a flawed Tetris game where pieces can move upward, exposing the unspecified allowed movement in requirement. The visual counterfactual is currently limited to the controllable game code generation; as GPTs outputs tend to be less predictable, we display static chat histories as illustrative examples for GPTs.
@@
We use OpenAI GPT-4o to power the interface, [^13] The state-of-the-art LLM at the time of writing: https://openai.com/index/hello-gpt-4o/. We used a temperature of 0.3 for code generation and 0.7 for other tasks. Refer to supplemental materials for our prompts. leveraging its interactive nature to adaptively provide feedback on the diverse set of possible user inputs. However, feedback generation does not rely on the model’s inherent reasoning capabilities but is anchored in our predefined reference requirements. Specifically, we always have GPT-4o to compare the current user requirements to the reference, and select the most critical defect to provide feedback on (roughly, mismatched requirements are considered most critical, then missing requirements, then ambiguous ones). Then, we have the LLM respond targetedly to the defect. For example, visual counterfactuals on game tasks are generated by having the LLM minimally edit reference Python code based on the identified incorrect requirement. To enhance interaction, we also implement features like highlighting special variables (e.g., board, keystrokes) in conversations and as hyperlinks for easy cross-referencing across the interface.
@@
We analyze the interface log data to evaluate the feedback quality for all types of feedback (Figure 4A, B, C). From our user study (Section 5), we collected a total of 635 interaction turns from Tetris (ranging from 11 to 73 turns per user) and 180 turns from Email Proofreader (ranging from 4 to 20 turns per user). From this, we randomly selected 10 different users and annotated conversations for 5 users per task, resulting in 151 Tetris turns, 66 Email Proofreader turns, and a total of 113 annotations per feedback type. For each LLM turn, we assess: (1) Is the feedback needed? (2) Is the feedback provided? (3) Is the provided feedback correct? Given the answers to the three questions, each feedback was categorized as correct, incorrect, irrelevant, missing, or not provided (Table 2). For example, irrelevant feedback are those that are (1) not needed but (2) provided. Note that the right feedback includes both correct feedback that was (1) needed, (2) provided, and (3) correct, and not provided feedback that was (1) not needed and (2) not provided.
@@
Description of the table. Feedback Source Right Feedback Wrong Feedback IRR ($\alpha$) Correct Not Provided Total Incorrect Irrelevant Missing Total Chat (e.g., Fig. 4A) 69.03% 20.35% 89.38% 10.62% 0.00% 0.00% 10.62% 0.86 Ref Example (Fig. 4B) 48.67% 38.05% 86.73% 1.77% 5.31% 6.19% 13.27% 0.88 LLM Output (Fig. 4C) 3.75% 86.25% 90.00% 0.00% 5.00% 5.00% 10.00% 0.68 All Feedback 44.44% 44.12% 88.56% 4.58% 3.27% 3.59% 11.44% 0.87